in the real numbers what is the solution of the equation 8^2x+1 = 4^1-x?
could i please have some work with this as well bcuz my answer key has c but i don't know how it got that so an explanation would be much appreciated. Thanks for ur kind generosity and help!
a.-1/3
b.-1/4
c.-1/8
d.0
e.1/7

Answer by lwsshak3(11628)

what is the solution of the equation 8^2x+1 = 4^1-x
**
8^(2x+1) = 4^(1-x)
8=2^3, 4=2^2
2^3^(2x+1) = 2^2^(1-x)
2^3(2x+1) = 2^2(1-x)
2^(6x+3)=2^(2-2x)
equate exponents with same base
6x+3=2-2x
8x=-1
x=-1/8
correct ans is c. above